# Dataset Distillation-based Hybrid Federated Learning on Non-IID Data

### v2

@@ VI-D Impact of distilled data size
The size of the distilled dataset significantly affects the amount of information retained from the local data of cluster members, which directly influences the performance of the global model. To investigate this impact, we evaluate the convergence behavior of HFLDD under varying levels of data heterogeneity on the CIFAR10 dataset, using different distilled data sizes. As shown in Fig. 7, increasing the distilled data size leads to notable improvements in both convergence speed and final test accuracy. This observation aligns with the theoretical expectation: a larger distilled dataset can better approximate the original data distribution, thereby reducing the distillation approximation error and improving the performance of the global model.
VI-E Impact of Network Architecture
To evaluate the generalizability and adaptability of our proposed framework across various network architectures, we conducted additional experiments using a diverse set of models on the CIFAR10 dataset, including MobileNetV2[54], ResNet18[55], and ShuffleNet[56]. These models cover a broad spectrum of architectural characteristics, such as lightweight mobile models and deeper residual networks.
Table V lists the test accuracies of different methods using the above three models under different levels of client heterogeneity. Overall, HFLDD exhibits stable and competitive performance across all model architectures and heterogeneity settings. Under severe non-IID settings, i.e., when $C=1$ and $C=2$, the test accuracies of HFLDD are largely higher than FedAvg and FedSeq, and mostly higher than that of DynaFed. Under the IID setting, i.e., when $C=10$, HFLDD remains competitive with the baseline methods, demonstrating strong adaptability across different model architectures.
TABLE V: Test accuracies of different methods using different models on CIFAR10
@@ VI-E Impact of Network Architecture
TABLE VI: The impact of $K$ on the test accuracy of HFLDD
@@ VI-F Impact of $K$ in K-Means clustering
In this part, we evaluate the impact of $K$ in the homogeneous clustering of K-Means clients on the performance of the proposed HFLDD. Table VI shows the test accuracies of HFLDD in the MNIST, CIFAR10 and CINIC10 datasets with different values of $K$. When $C=1$, on the CIFAR10 and CINIC10 dataset, the test accuracy is the highest when $K=10$, and shows a noticeable decline as $K$ decreases or increases. However, in the MNIST dataset, the test accuracy remains relatively unchanged as $K$ varies. This is because the MNIST dataset is relatively simple, making its data patterns and features easy to capture and understand. In this scenario, even if an inappropriate value of $K$ is selected, HFLDD still can easily learn the features in the MNIST dataset.
When $C=2$ and $C=10$, heterogeneity is mitigated and there is partial overlap in label classes between different clients, leading to a certain degree of overlap in the feature space. The test accuracy on the three datasets remains relatively stable as the value of $K$ changes. From this, we observe that when the heterogeneity of the data is not particularly severe, HFLDD exhibits robustness to the value of $K$.
In addition to model accuracy, we further examine the effect of $K$ on the convergence rate. As illustrated in Fig. 8. From the figure, we observe that the number of clusters has a relatively small effect on the convergence speed.
(a) $C=1$
(b) $C=2$
(c) $C=10$
Figure 8: Learning curves of the global model under different homogeneous cluster numbers across varying data heterogeneity on CIFAR10.
@@ VII Discussion
It is important to note that the actual scenario differs from the example provided above, the data heterogeneity among clients’ data types is highly complex. After sampling and reorganizing homogeneous clusters, it is highly possible to have a limited number of clients in some heterogeneous clusters. There may even be many isolated nodes considered as a single heterogeneous cluster. This leads to an uneven number of clients among clusters. In decentralized settings, if these uneven clusters are linked, as in Ref. [32], to form a ring-shaped topology for collaborative training, it will lead to a significant decrease in convergence speed and test accuracy. The most intuitive approach to address this issue is to directly remove heterogeneous clusters with isolated or a very small number of clients from training. However, this often results in worse training performance, which is unable to fully utilize the data in each client.
@@ VIII Conclusion
In this paper, we propose a communication-efficient HFL framework that addresses the challenge of distributed learning with non-IID data by leveraging dataset distillation. Our approach constructs heterogeneous clusters based on the knowledge of the labels of all clients. Within each heterogeneous cluster, members transmit distilled data to a cluster head. This process aims to generate approximately IID datasets among clusters. Extensive experiments demonstrate that our proposed HFLDD outperforms the considered baseline methods and can significantly reduce the impact of non-IID data on training performance with a much smaller communication cost.

### v3

@@ VI-D Impact of distilled data size
The size of the distilled data significantly affects the amount of information retained from the local data of cluster members, which directly influences the performance of the global model. To investigate this impact, we evaluate the convergence behavior of HFLDD under varying levels of data heterogeneity on CIFAR10, using different distilled data sizes. As shown in Fig. 7, increasing the distilled data size leads to notable improvements in both convergence speed and final test accuracy. This observation aligns with the theoretical expectation: a larger distilled data can better approximate the original data distribution, thereby reducing the distillation approximation error and improving the performance of the global model.
TABLE VI: Test accuracies with different model architectures on CIFAR10
@@
VI-E Impact of Network Architecture
To evaluate the generalizability and adaptability of our proposed framework across various network architectures, we conducted additional experiments using a diverse set of models on CIFAR10, including MobileNetV2[sandler2018mobilenetv2], ResNet18[he2016deep], and ShuffleNet[zhang2018shufflenet]. These models cover a broad spectrum of architectural characteristics, such as lightweight mobile models and deeper residual networks.
Table VI lists the test accuracies of different methods using the above three models under different levels of client heterogeneity. Overall, HFLDD exhibits stable and competitive performance across all model architectures and heterogeneity settings. Under severe non-IID settings, i.e., when $C=1$ and $C=2$, the test accuracies of HFLDD are largely higher than FedAvg and FedSeq, and mostly higher than that of DynaFed. Under the IID setting, i.e., when $C=10$, HFLDD remains competitive with the baseline methods, demonstrating strong adaptability across different model architectures.
TABLE VII: The impact of $K$ on the test accuracy of HFLDD
@@ VI-F Impact of $K$ in K-Means clustering
In this part, we evaluate the impact of $K$ on the homogeneous clustering of K-Means clients on the performance of HFLDD. Table VII shows the test accuracies of HFLDD in MNIST, CIFAR10, and CINIC10 with different values of $K$. When $C=1$, on CIFAR10 and CINIC10, HFLDD achieves the highest test accuracy when $K=10$, and shows a noticeable decrease as $K$ decreases or increases. However, in MNIST, the test accuracy remains relatively unchanged as $K$ varies. This is because MNIST is a relatively simple dataset, its data patterns and features are easy to capture and understand. In this scenario, even if an inappropriate value of $K$ is selected, HFLDD can still easily learn the features in the MNIST dataset.
When $C=2$ and $C=10$, there is overlap in the label classes between different clients, leading to a certain degree of overlap in the feature space. The test accuracies on the three datasets remain relatively stable as the value of $K$ changes. We can see that when data heterogeneity is not particularly severe, HFLDD exhibits robustness to the value of $K$.
@@ VII Discussion
It is important to note that the actual scenario differs from the example provided above, the data heterogeneity among clients’ data types is highly complex. After sampling and reorganizing homogeneous clusters, it is highly possible to have a limited number of clients in some heterogeneous clusters. There may even be many isolated nodes considered as a single heterogeneous cluster. This leads to an uneven number of clients among clusters. In decentralized settings, if these uneven clusters are linked, as in Ref. [abebe2023optimizing], to form a ring-shaped topology for collaborative training, it will lead to a significant decrease in convergence speed and test accuracy. The most intuitive approach to address this issue is to directly remove heterogeneous clusters with isolated or a very small number of clients from training. However, this often results in worse training performance, which is unable to fully utilize the data in each client.
@@ VIII Conclusion
In this paper, we propose a communication-efficient HFL framework that addresses the challenge of distributed learning with non-IID data by leveraging dataset distillation. Our approach constructs heterogeneous clusters based on the knowledge of the labels of all clients. Within each heterogeneous cluster, members transmit distilled data to a cluster head. This process generates approximately IID datasets among clusters. Extensive experiments demonstrate that our proposed HFLDD outperforms the considered baseline methods and can significantly reduce the impact of non-IID data on training performance with a much smaller communication cost.
While the proposed HFLDD framework demonstrates significant advantages in handling non-IID data, it has several limitations that need further investigation. First, HFLDD relies on dataset distillation to compress local data into distilled data, which inevitably introduces information loss, particularly under highly imbalanced or long-tailed label distributions. Second, although distilled data are generally considered to offer privacy protection, potential risks may still exist in highly sensitive applications. Third, the current clustering mechanism does not consider system heterogeneity and communication instability, which may affect cluster formation and overall training performance.
We plan to address these limitations in future work by developing lightweight dataset distillation algorithms that are better suited for long-tailed data distributions to reduce information loss and computational overhead, investigating privacy-enhancing techniques for distilled data, and extending our framework by incorporating adaptive clustering strategies and more resilient communication mechanisms that can support dynamic participation and improve resilience to communication instability.